MKTEXPK(1) MKTEXPK(1) NAME mktexpk−createaPKfileforafont SYNOPSIS mktexpk[options]font[redirect] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualKpathsea:Alibraryforpathsearching. mktexpkisusedtogenerateapkfilefromtheMETAFONTsourcefilesortype1sourcesforfont, ifpossible. Ifredirectisgiven,itmustbestringoftheform>&nwherenisthenumberofthefiledescriptor whichistoreceive,insteadofstandardoutput,thenameofthegeneratedfile. mktexpkistypicallycalledbyotherprograms,ratherthanfromthecommandline. OPTIONS mktexpkacceptsthefollowingoptions: --bdpinum Setthebasedeviceresolution. --dpinum Generatethefontatresolutionnum. --helpPrinthelpmessageandexit. --magmag TheMETAFONTmagnificationthatcorrespondstothecombinationofbdpianddpithat hasbeenspecified. --mfmodemode TheMETAFONTmodetouse. --destdirdir Adirectoryname.Ifthedirectoryisabsolute,itisusedas-is.Otherwise,itisappended totherootdestinationdirectorysetinthescript. --version Printversioninformationandexit. BUGS Ifthemodeisnotspecified,mktexpktriestodetermineitfrombdpi.Themannerinwhichthis isdoneisprimitive.Sincebpdidoesn’tprovidesufficientinformationtouniquelydeterminea modeitisstronglyrecommendedthatitbespecifiedexplicitly. Ifthemodeandbpdidon’tmatch,themodeisdiscarded. Thedpiparameterisredundant,butisnotcheckedforconsistency.Worse,theprogramwillfail ifthewrongdpiisspecified. SEEALSO mktexmf(1),mktextfm(1). Kpathsea7.32x 4January1998 1